# RestoreGrad: Signal Restoration Using Conditional Denoising Diffusion Models with Jointly Learned Prior

### v1

@@ 6 Conclusion
Combining (22) and (23), we have obtained the --ELBO of Proposition 3.2.
@@ C.2.1 Dataset
We used the VoiceBank+DEMAND dataset (Valentini-Botinhao et al., 2016) with the same experimental setup as in previous work (Pascual et al., 2017; Phan et al., 2020; Strauss & Edler, 2021; Lu et al., 2022) to perform a direct comparison. The clean speech and noise recordings were provided from the VoiceBank corpus (Veaux et al., 2013) and the Diverse Environments Multichannel Acoustic Noise Database (DEMAND) (Thiemann et al., 2013), respectively, each recorded with sampling rate of 48kHz. Noisy speech inputs used for training were composed by mixing the two datasets with four signal-to-noise ratio (SNR) settings from {0, 5, 10, 15} dB, using 10 types of noise (2 artificially generated + 8 real recorded from the DEMAND dataset) and 28 speakers from the Voice Bank corpus. The test set inputs were made with four SNR settings different from the training set, i.e., {2.5, 7.5, 12.5, 17.5} dB, using the remaining 5 noise types from DEMAND and 2 speakers from the VoiceBank corpus. There are totally 11527 utterances for training and 824 for testing. Note that the speaker and noise classes were uniquely selected for the training and test sets. The dataset is publicly available at: https://datashare.ed.ac.uk/handle/10283/2826. In our experiments, the audio steams were resampled to 16kHz sampling rate.
@@ C.2.3 Optimization and Inference
We used the same configurations of CDiffuSE (Base) (Lu et al., 2022) for optimizing all the models, where the batch size was 16, the Adam optimizer was used with a learning rate of $2\times 10^{-4}$, and the diffusion steps $T=50$ with linearly spaced $\beta_{t}\in[10^{-4},0.035]$. For RestoreGrad, we imposed the minimum standard deviation $\sigma_{\text{min}}=0.1$ by adding it to the output of the Prior Net and Posterior Net to ensure stability during training. The fast sampling scheme in Kong et al. (2021) was used in the reverse processes with $S=6$ and the inference schedule $\beta^{\text{infer}}_{t}=[10^{-4},10^{-3},0.01,0.05,0.2,0.35]$. The models were trained on one NVIDIA Tesla V100 GPU of 32 GB CUDA memory and finished training for 96 epochs in 1 day.
@@ C.3.1 Datasets
Snow100K (Liu et al., 2018): a dataset for evaluation of image desnowing models. We used the test set for evaluation, which consist of 50,000 samples. The images are split into approximately equal sizes of three Snow100K-S/M/L sub-test sets (16,611/16,588/16,801), indicating the synthetic snow strength imposed via snowflake sizes (light/mid/heavy). The dataset can be downloaded from: https://sites.google.com/view/yunfuliu/desnownet.
@@ C.3.3 Optimization and Inference
We used the same configurations of Özdenizci & Legenstein (2023) for optimizing all the models, except the batch size was 4 instead of 16 due to GPU memory limitation. The Adam optimizer with a fixed learning rate of $2\times 10^{-5}$ was used for training models without weight decay, and an exponential moving average with a weight of 0.999 was applied during parameter updates. The number of diffusion steps was $T=1000$ and the noise schedule was $\beta_{t}\in[10^{-4},0.02]$, linearly spaced. For inference, we used $S=10$ sampling timesteps for each model that we trained on our own. We did not use the deterministic implicit sampling scheme as in Özdenizci & Legenstein (2023) for our RestoreGrad-based DDPM models as we found using the normal stochastic sampling scheme actually works better. The models were trained on 2 NVIDIA Tesla V100 GPU of 32 GB CUDA memory and finished training for 9,261 epochs on the RainDrop dataset in 12 days and 887 epochs on the AllWeather dataset in 21 days.
@@ C.3.4 Evaluation Metrics
SSIM: a non-linear full-reference metric compares the luminance, contrast and structure of the original and degraded image (Wang et al., 2004). It provides a value from 0 to 1, the closer the score is to 1, the more similar the degraded image is to the reference image. We followed Özdenizci & Legenstein (2023) to compute SSIM based on the luminance channel Y of the YCbCr color space. We used the implementation form https://github.com/JingyunLiang/SwinIR for SSIM calculation.
Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) and Fréchet Inception Distance (FID) (Heusel et al., 2017): to provide better quantification of perceptual quality over the traditional distortion measures of PSNR and SSIM (Blau & Michaeli, 2018; Freirich et al., 2021). For the LPIPS we used the implementation from https://github.com/richzhang/PerceptualSimilarity, and for FID we used the implementation from https://github.com/chaofengc/IQA-PyTorch. In both metrics, a lower score indicates better perceptual quality of the restored image.
@@ C.4.1 Datasets
CHiME-3 (Barker et al., 2017): is a 6-channel microphone recording of talkers speaking in a noisy environment, sampled at 16 kHz. It consists of 7138 and 1320 simulated utterances for training and testing, respectively, which are generated by artificially mixing clean speech data with noisy backgrounds of four types, i.e. cafe, bus, street, and pedestrian area. In this paper, we only take the 5-th channel recordings for the experiments. The dataset information can be found at: https://www.chimechallenge.org/challenges/chime3/data.
@@ Appendix D Additional Experimental Results
D.1 Generalization to Out-of-Distribution (OOD) and Realistic Data
We have so far evaluated the models on in-domain scenarios with synthetic noisy data where RestoreGrad has shown substantial improvements. A natural question is that if the demonstrated improvements have actually come at the expense of the model’s generalizability to unseen or realistic data. To address the concern, we evaluate the IR models on two additional datasets from Quan et al. (2021); Liu et al. (2018) that consist of real-world images, using the reference-free Natural Image Quality Evaluator (NIQE) metric (Mittal et al., 2012) (a lower score indicates better quality). In Table D.1 we see that RestoreGrad is able to perform on par with or better than WeatherDiff and the non-generative TransWeather model. For OOD testing, we evaluate the SE models on the CHiME-3 dataset (Barker et al., 2017) unseen during model training. Table D.1 compares RestoreGrad with CDiffuSE that was also trained for 96 epochs, DOSE (Tai et al., 2023a), and two discriminative SE models. We can see that RestoreGrad is able to perform equally well as the CDiffuSE while outperforming DOSE and other non-generative SE models (Demucs (Defossez et al., 2020), WaveCRN (Hsieh et al., 2020)). The results in both tables show that RestoreGrad is capable of improving in-domain performance while maintaining desirable generalization capabilities of generative models.
Table 5: Evaluation on realistic image datasets of IR models trained on synthetic images of AllWeather training set.
{NiceTabular}
lccc Methods Gen. RainDS-Real Snow-Real
NIQE $\downarrow$ NIQE $\downarrow$
TransWeather (Valanarasu et al., 2022) N 4.005 3.161
WeatherDiff (Özdenizci & Legenstein, 2023) Y 3.050 2.985
+ RestoreGrad (ours) Y 2.556 3.015
{NiceTabular}
[…]
Unprocessed - 1.27 2.61 1.93 1.88 7.51
Demucs (Defossez et al., 2020) N 1.38 2.50 2.08 1.88 -
WaveCRN (Hsieh et al., 2020) N 1.43 2.53 2.03 1.91 -
DOSE (Tai et al., 2023a) Y 1.52 2.71 2.15 2.06 -
CDiffuSE (Lu et al., 2022) Y 1.55 2.87 2.09 2.15 7.67
+ RestoreGrad (ours) Y 1.54 2.88 2.14 2.16 8.45
D.2 Additional Results on SE
@@
Figure 9: Effect of using reduced numbers of sampling steps in inference on the SE performance, in terms of PESQ. RestoreGrad demstrates strongest endurance to the reduction in reverse sampling steps for inference.
@@
Visualizing the Learned Prior. It would be interesting to see how the latent noise prior that has been learned by RestoreGrad looks like and how it compares with that of the PriorGrad. In Figure 10 we present an example of a randomly chosen noisy speech waveform and the corresponding latent noise $\boldsymbol{\Sigma}_{y}=\mbox{diag}\{\boldsymbol{\sigma}^{2}_{y}\}$ of PriorGrad and that of RestoreGrad (with $(\eta,\lambda)=(0.1,0.5)$ for (11)). It can be seen that the variances of the pre-defined (PriorGrad) and learned (RestoreGrad) latent noise distributions are actually quite different, though both show the trend of following the variation of the conditioner signal level. This trend indicates that both latent distributions aim to better approximate the true signal distribution in a more informative manner for improved efficiency, as against the standard Gaussian prior used in the original DDPM. Note that in the RestoreGrad training, we have chosen a proper KL weight $\lambda$ so that the Prior Net distribution matches the Posterior Net distribution reasonably well without harming the reconstruction ability of the DDPM model. On the other hand, using a too large $\lambda$ might lead to a collapsed latent space as the optimization could have put too much emphasis on matching the prior and posterior distribution, discarding the contribution of the reconstruction loss term. In contrast, using a too small $\lambda$ might result in large discrepancy between the learned prior and posterior distributions, as also illustrated in Figure 10. Empirically, we found a naive choice of 1 works reasonably well and also for similar values, e.g., 0.5, 10, etc., as similarly observed in the VAE-type model of Kohl et al. (2018).
@@
Evaluation Using Automatic Speech Recognition (ASR): Following Benita et al. (2024) who perform evaluation of diffusion-based speech generation using ASR, we evaluate the SE model as a front-end denoiser for ASR under noisy environments. To this end, we pre-process the noisy VoiceBand+DEMAND test data samples through the well-trained SE model and feed the denoised audio separately to two pre-trained ASR engines taken from the NVIDIA NeMo toolkit [^1] https://github.com/NVIDIA/NeMo: Conformer-transducer-large (Gulati et al., 2020) and Citrinet-1024 (Majumdar et al., 2021). We report the word error rate (WER) and character error rate (CER) for each ASR engine outcome, where the lower WER / CER indicate better performance. The results are presented in Table D.2 with all the SE models trained after 96 epochs, inferred using 6 steps. It is interesting to see that CDiffuSE and PriorGrad actually lead to worse performance than the unprocessed speech case for Citrinet ASR. Our RestoreGrad is able to achieve the lowest WER and CER for both ASR models, demonstrating its efficacy for enhancing machine learning capabilities under noisy environments.
Table 9: Following Benita et al. (2024) who perform evaluation of diffusion-based speech generation using ASR, we evaluate SE models on two ASR engines (Conformer, Citrinet) for the VoiceBand+DEMAND test set. The results further confirm the superiority of RestoreGrad over the baseline and PriorGrad.
@@ D.3 Additional Results on IR
RainDropDiff Özdenizci & Legenstein (2023) 32.29 0.9422
@@
Restoration Performance vs. $\eta$ and $\lambda$: We also study the IR performance of the RestoreGrad models trained across various combinations of $\eta$ and $\lambda$ in Table D.3, where the models were trained and tested on the RainDrop dataset. The results show that RestoreGrad works effectively for a wide range of $\eta$ and $\lambda$ values as outperforming the baseline DDPM model, RainDropDiff from Özdenizci & Legenstein (2023), which utilizes the standard Gaussian prior for the diffusion process.
Table 11: RestoreGrad performance for various $\eta$ and $\lambda$, where the models were trained for 9,261 epochs and tested with $S=10$ sampling steps on the RaindDrop dataset (Qian et al., 2018). The baseline RainDropDiff model results reported in the original paper of Özdenizci & Legenstein (2023) (which was trained for 37,042 epochs, 4 times more than our RestoreGrad models) are also presented here for comparison purposes.
@@
More Image Restoration Examples: We provide more examples in Figures 13, 14, 15, 16, 17 for comparing our RestoreGrad with the baseline DDPM approaches (i.e., RainDropDiff, WeatherDiff) of Özdenizci & Legenstein (2023). The restored images of WeatherDiff were obtained by using the trained model weights provided by Özdenizci & Legenstein (2023) at https://github.com/IGITUGraz/WeatherDiffusion. We also provide examples of image deblurring in Figures 18 and 19.
Figure 13: Restored images by RainDropDiff (Özdenizci & Legenstein, 2023) and RestoreGrad (ours) for a test sample from the RainDrop dataset.
Figure 14: Image restoration examples using a test image taken from the Snow100K-L test set.
Figure 15: Image restoration examples using a test image taken from the Outdoor-Rain test set.
Figure 16: Image restoration examples using a test image taken from the RainDrop test set.
Figure 17: Image restoration examples using a test image taken from the RainDrop test set.

### v3

@@ 6 Conclusion
Combining (22) and (23), we have obtained the --ELBO $\mathcal{L}(\theta,\phi,\psi)$ of Proposition 3.2.
@@ C.2.1 Dataset
We used the VoiceBank+DEMAND dataset (Valentini-Botinhao et al., 2016) with the same experimental setup as in previous work (Pascual et al., 2017; Phan et al., 2020; Strauss & Edler, 2021; Lu et al., 2022) to perform a direct comparison. The clean speech and noise recordings were provided from the VoiceBank corpus (Veaux et al., 2013) and the Diverse Environments Multichannel Acoustic Noise Database (DEMAND) (Thiemann et al., 2013), respectively, each recorded with sampling rate of 48kHz. Noisy speech inputs used for training were composed by mixing the two datasets with four signal-to-noise ratio (SNR) settings from {0, 5, 10, 15} dB, using 10 types of noise (2 artificially generated + 8 real recorded from the DEMAND dataset) and 28 speakers from the Voice Bank corpus. The test set inputs were made with four SNR settings different from the training set, i.e., {2.5, 7.5, 12.5, 17.5} dB, using the remaining 5 noise types from DEMAND and 2 speakers from the VoiceBank corpus. There are totally 11,527 utterances for training and 824 for testing. Note that the speaker and noise classes were uniquely selected for the training and test sets. The dataset is publicly available at: https://datashare.ed.ac.uk/handle/10283/2826. In our experiments, the audio steams were resampled to 16kHz sampling rate.
@@ C.2.3 Optimization and Inference
We used the same configurations of CDiffuSE (Base) (Lu et al., 2022) for optimizing all the models, where the batch size was 16, the Adam optimizer was used with a learning rate of $2\times 10^{-4}$, and the diffusion steps $T=50$ with linearly spaced $\beta_{t}\in[10^{-4},0.035]$. For RestoreGrad, we imposed the minimum standard deviation $\sigma_{\text{min}}=0.1$ by adding it to the output of the Prior Net and Posterior Net to ensure stability during training. The fast sampling scheme in Kong et al. (2021) was used in the reverse processes with 6-step inference schedule $\beta^{\text{infer}}_{t}=[10^{-4},10^{-3},0.01,0.05,0.2,0.35]$. The models were trained on one NVIDIA Tesla V100 GPU of 32 GB CUDA memory and finished training for 96 epochs in 1 day.
@@ C.3.1 Datasets
Snow100K (Liu et al., 2018): a dataset for evaluation of image desnowing models. The images are split into approximately equal sizes of three Snow100K-S/M/L sub-test sets (with 16,611/16,588/16,801 samples, respectively), indicating the synthetic snow strength imposed via snowflake sizes (light/mid/heavy). We used the Snow100K-L sub-test set for evaluation. The dataset can be downloaded from: https://sites.google.com/view/yunfuliu/desnownet.
@@ C.3.3 Optimization and Inference
We used the same configurations of Özdenizci & Legenstein (2023) for optimizing all the models, except the batch size was 4 instead of 16 due to GPU memory limitation. The Adam optimizer with a fixed learning rate of $2\times 10^{-5}$ was used for training models without weight decay, and an exponential moving average with a weight of 0.999 was applied during parameter updates. The number of diffusion steps was $T=1000$ and the noise schedule was $\beta_{t}\in[10^{-4},0.02]$, linearly spaced. For inference, we used the sampling scheme with 10 timesteps for each model that we trained on our own. We did not use the deterministic implicit sampling scheme as in Özdenizci & Legenstein (2023) for our RestoreGrad-based DDPM models as we found using the normal stochastic sampling scheme actually works better. The models were trained on 2 NVIDIA Tesla V100 GPU of 32 GB CUDA memory and finished training for 9,261 epochs on the RainDrop dataset in 12 days and 887 epochs on the AllWeather dataset in 21 days.
@@ C.3.4 Evaluation Metrics
SSIM: a non-linear full-reference metric compares the luminance, contrast and structure of the original and degraded image (Wang et al., 2004). It provides a value from 0 to 1, the closer the score is to 1, the more similar the degraded image is to the reference image. We followed Özdenizci & Legenstein (2023) to compute SSIM based on the luminance channel Y of the YCbCr color space. We used the implementation form https://github.com/JingyunLiang/SwinIR for SSIM.
LPIPS (Zhang et al., 2018) and FID (Heusel et al., 2017): to provide better quantification of perceptual quality over the traditional distortion measures of PSNR and SSIM (Blau & Michaeli, 2018; Freirich et al., 2021). For the LPIPS we used the implementation from https://github.com/richzhang/PerceptualSimilarity, and for FID we used the implementation from https://github.com/chaofengc/IQA-PyTorch. In both metrics, a lower score indicates better perceptual quality of the restored image.
@@ C.4.1 Datasets
CHiME-3 (Barker et al., 2017): is a 6-channel microphone recording of talkers speaking in a noisy environment, sampled at 16 kHz. It consists of 7,138 and 1,320 simulated utterances for training and testing, respectively, which are generated by artificially mixing clean speech data with noisy backgrounds of four types, i.e. cafe, bus, street, and pedestrian area. In this paper, we only take the 5-th channel recordings for the experiments. The dataset information can be found at: https://www.chimechallenge.org/challenges/chime3/data.
@@ Appendix D Additional Experimental Results
D.1 Additional Results on SE
@@
Figure 9: Effect of using reduced numbers of sampling steps in inference on the SE performance, in terms of PESQ. RestoreGrad demonstrates the strongest resistance to the reduction in reverse sampling steps for inference.
@@
Visualizing the Learned Prior. It would be interesting to see how the latent noise prior that has been learned by RestoreGrad looks like and how it compares with that of the PriorGrad. In Figure 10 we present an example of a randomly chosen noisy speech waveform and the corresponding latent noise $\boldsymbol{\Sigma}_{y}=\mbox{diag}\{\boldsymbol{\sigma}^{2}_{y}\}$ of PriorGrad and that of RestoreGrad (with $(\eta,\lambda)=(0.1,0.5)$ for (11)). It can be seen that the variances of the pre-defined (PriorGrad) and learned (RestoreGrad) latent noise distributions are actually quite different, though both show the trend of following the variation of the conditioner signal level. This trend indicates that both latent distributions aim to better approximate the true signal distribution in a more informative manner for improved efficiency, as against the standard Gaussian prior used in the original DDPM. Note that in the RestoreGrad training, we have chosen a proper KL weight $\lambda$ so that the Prior Net distribution matches the Posterior Net distribution reasonably well without harming the reconstruction performance of the DDPM model. On the other hand, using a too large $\lambda$ might lead to a collapsed latent space as the optimization could have put too much emphasis on matching the prior and posterior distribution, discarding the contribution of the reconstruction loss term. In contrast, using a too small $\lambda$ might result in large discrepancy between the learned prior and posterior distributions, as also illustrated in Figure 10. Empirically, we found a naive choice of 1 works reasonably well and also for similar values, e.g., 0.5, 10, etc., as similarly observed in the VAE-type model of Kohl et al. (2018).
@@
Evaluation Using Automatic Speech Recognition (ASR): Following Benita et al. (2024) who perform evaluation of diffusion-based speech generation using ASR, we evaluate the SE model as a front-end denoiser for ASR under noisy environments. To this end, we pre-process the noisy VoiceBand+DEMAND test data samples through the well-trained SE model and feed the denoised audio separately to two pre-trained ASR engines taken from the NVIDIA NeMo toolkit [^1] https://github.com/NVIDIA/NeMo: Conformer-transducer-large (Gulati et al., 2020) and Citrinet-1024 (Majumdar et al., 2021). We report the word error rate (WER) and character error rate (CER) for each ASR engine outcome, where the lower WER / CER indicate better SE performance. The results are presented in Table D.1 with all the SE models trained after 96 epochs, inferred using 6 steps. It is interesting to see that CDiffuSE and PriorGrad actually lead to worse performance than the unprocessed speech case for Citrinet ASR. Our RestoreGrad is able to achieve the lowest WER and CER for both ASR models, demonstrating its efficacy for enhancing machine learning capabilities under noisy environments.
Table 9: Following Benita et al. (2024) who perform evaluation of diffusion-based speech generation using ASR, we evaluate SE models on two ASR engines (Conformer, Citrinet) for the VoiceBand+DEMAND test set. The results further confirm the superiority of RestoreGrad over the baseline DDPM (CDiffuSE) and PriorGrad.
@@ D.2 Additional Results on IR
RDDM (Liu et al., 2024) 32.51 0.9563
RainDropDiff (Özdenizci & Legenstein, 2023) 32.29 0.9422
@@
Restoration Performance vs. $\eta$ and $\lambda$: We also study the IR performance of the RestoreGrad models trained across various combinations of $\eta$ and $\lambda$ in Table D.2, where the models were trained and tested on the RainDrop dataset. The results show that RestoreGrad works effectively for a wide range of $\eta$ and $\lambda$ values to outperform the baseline DDPM model, RainDropDiff from Özdenizci & Legenstein (2023), which utilizes the standard Gaussian prior for the diffusion process.
Table 11: RestoreGrad performance for various $\eta$ and $\lambda$, where the models were trained on the RaindDrop training set (Qian et al., 2018) for 9,261 epochs and evaluated on the test set. The baseline RainDropDiff model scores reported in the original paper of Özdenizci & Legenstein (2023) (which was trained for 37,042 epochs, 4 times more than our RestoreGrad models) are also presented here for comparison purposes.
@@
Images Generated from Different Prior Noise Samples: In Figure 13, we present example images generated for the same conditioner $\mathbf{y}$ using different random seed values (1, 10, and 20) when sampling the latent noise $\boldsymbol{\epsilon}$, for both the baseline DDPM (RainDropDiff) and our RestoreGrad. In the figure, although it is challenging to perceive the difference between the results of different seed values simply by inspecting the images visually, the quality (PSNR and SSIM scores) of the images produced by RestoreGrad is consistently better than the baseline DDPM among the different seed values used, further demonstrating the superiority of our method in restoring higher fidelity signals.
Figure 13: Example restored images from different prior noise samples (by using different random seed values) for the baseline DDPM (RainDropDiff) and our approach (RestoreGrad).
Signal Quality and Encoder Complexity Trade-Offs: Similar to the computational complexity analysis provided in Table 2 for the SE task, we also present the results for the IR task on the RainDrop dataset in Table 12. From the table, we again observe that the performance of our approach slightly improves with the use of a larger encoder, while also with increasing complexity. However, the computational overhead of our approach, i.e., the additional complexity due to the encoder module(s), is again relatively smaller compared to the adopted DDPM module (i.e., RainDropDiff (Özdenizci & Legenstein, 2023)), in terms of processing latency ($<$1.3% of DDPM) and memory usage ($<$30% of DDPM).
Table 12: IR comparison of RestoreGrad models (on RainDrop dataset) using encoder modules of different sizes and the corresponding latency and GPU memory usage (measured on one NVIDIA Tesla V100 GPU) presented as the ratio of encoder to DDPM.
[…]
Base (0.27M) 32.65 0.9414 0.9% 16%
Large (1.9M) 32.77 0.9444 1.3% 30%
Visualization of Diffusion Processes: We present a qualitative analysis of the diffusion processes by showing a sequence of images from $t=0$ to $t=T$ for the baseline conditional diffusion model (WeatherDiff) and with the proposed method (RestoreGrad) in Figure 14. It can be seen that, at a given timestep $t$, the generated image by using RestoreGrad is cleaner than that of WeatherDiff, indicating a better diffusion trajectory developed by using our approach.
Figure 14: Visualization of diffusion processes for the baseline DDPM (WeatherDiff) and our method (RestoreGrad).
@@
More Image Restoration Examples: We provide more examples in Figures 15, 16, 17, 18, 19, for comparing our RestoreGrad with the baseline DDPM approaches (i.e., RainDropDiff, WeatherDiff) of Özdenizci & Legenstein (2023). The restored image of RainDropDiff in Figure 15 was obtained by using the model weights trained by ourselves. The restored images of WeatherDiff in Figures 16, 17, 18, 19 were obtained by using the trained model weights provided by Özdenizci & Legenstein (2023) at https://github.com/IGITUGraz/WeatherDiffusion. We also provide examples of image deblurring in Figures 20 and 21.
Figure 15: Restored images by RainDropDiff (Özdenizci & Legenstein, 2023) and RestoreGrad (ours) for a test sample from the RainDrop test set.
Figure 16: Restored images by WeatherDiff (Özdenizci & Legenstein, 2023) and RestoreGrad (ours) for a test sample from the Outdoor-Rain test set.
Figure 17: Restored images by WeatherDiff (Özdenizci & Legenstein, 2023) and RestoreGrad (ours) for a test sample from the Snow100K-L test set.
Figure 18: Restored images by WeatherDiff (Özdenizci & Legenstein, 2023) and RestoreGrad (ours) for a test sample from the RainDrop test set.
Figure 19: Restored images by WeatherDiff (Özdenizci & Legenstein, 2023) and RestoreGrad (ours) for a test sample from the RainDrop test set.